# Tailoring Dynamical Codes for Biased Noise: The X3Z3 Floquet Code

### v1

@@ V Results and Discussion
Figure 6: (a,b) Thresholds ($p_{\mathrm{th}}$) and (c,d) sub-threshold logical failure probability $p_{L}$ of different Floquet codes as a function of noise bias $\eta$. Codes studied are X3Z3 (blue), CSS (green), P6 (red) and XYZ2 (orange). Results are computed for two different noise models: (a,c) code-capacity and (b,d) SDEM3. (a) Inset: Zoom-in threshold plots for P6 (red) and XYZ2 (orange) codes. Lower panels: Results are calculated for code distance (c) Code size $12\times 18$ ($d=12$) with 18 QEC rounds with $p=0.72\%$ and (d) Code size $24\times 36$ ($d_{\mathrm{eff}}=12$) with 36 QEC rounds with $p=0.5\%$. Each data point is averaged over $10^{6}-10^{9}$ shots. For better visualisations, we fit all curves with quadratic splines.
@@
Figure 7: Exponential suppression of the sub-threshold logical failure probabilities $p_{L}$ with respect to $d_{\mathrm{eff}}$ for different Floquet codes: XYZ2 (orange), P6 (red), X3Z3 (blue), and CSS (green). The results are calculated for different noise models: (a,b) Code-capacity and (c,d) SDEM3 noise models, with a physical error rate $p=0.55\%$. They are also computed using different bias strengths: (a,c) $\eta=1$ and (b,d) $\eta=99$. All curves can be fitted to an exponential decay function $f\propto\mathrm{exp}(-\gamma d)$ or $f\propto\mathrm{exp}(-\gamma d_{\mathrm{eff}})$ where $\gamma$ depends on the bias strength $\eta$ and is an increasing function of $(p_{\mathrm{th}}-p)$. Each data point is averaged over $10^{6}-10^{9}$ shots.
@@
The upper panels of Fig. 6 show the thresholds of all codes for various levels of noise bias $\eta$ calculated for (a) code-capacity and (b) SDEM3 noise models. Each threshold is obtained from the intersection of the logical failure probabilities $p_{L}$ vs physical error rate $p$ curves of different code distances $d_{\mathrm{eff}}=d/2=6,8,10$ and 12. The data of $p_{L}$ as a function of $p$ for each code and various noise-bias strengths, used to obtain the thresholds, are presented in Figs. 11-14 of Appendix C.
@@
Since the code-capacity noise model which considers only single-qubit noise is a more benign model than the SDEM3 noise model which also includes two-qubit and measurement errors, the code performance calculated under the code-capacity noise is better than that of the SDEM3 noise. For both noise models, as shown in Figs. 6(a) and 6(b), the performance of the X3Z3 code becomes increasingly better than those of all other tested codes as the noise bias increases. In particular, as the noise changes from fully depolarising to a pure dephasing type, the X3Z3 Floquet code’s threshold increases from $\approx 1.13\%$ to $\approx 3.09\%$ for the code-capacity noise and from $\approx 0.76\%$ to $\approx 1.08\%$ for the SDEM3 noise model. The threshold therefore increases by a factor of 2.7 and a factor of 1.4 for the code capacity and SDEM3 noise models, respectively. We also investigate the performance of the X3Z3 Floquet code with a twisted periodic boundary condition where we find that the threshold at the infinite bias can reach up to $\approx 4.4\%$ (see Appendix F). This increased in the threshold of the twisted code is due to the fact that the pure $Z$-type logical operators has a length that scales quadratically with the distance of the untwisted code, thereby increases its code distance. We note, however, that the advantage of the twisted over the untwisted code only holds for the noise regime which is close to the infinite bias. As shown in Figs. 6(a) and 6(b), while the CSS Floquet code has the same threshold as the X3Z3 code when the noise is fully depolarising, its threshold decreases with increasing noise bias. This decrease is due to the fact that the CSS Floquet code has pure $X$ and pure $Z$ detectors: in the presence of biased noise, half of the detectors which are of the same type of the dominant error will become less useful in detecting the biased errors.
@@
Figures. 6(a) and 6(b) show that the thresholds for the honeycomb codes have only minor improvements as the noise bias increases. For the fully depolarising code-capacity noise, all codes have the same threshold, whereas for the SDEM3 depolarising noise the honeycomb code thresholds are about 1.3 times smaller than the CSS and X3Z3 Floquet code thresholds. This is partly explained by noting that the SDEM3 noise model contains measurement errors which give rise to hyperedges in the decoding hypergraphs of the honeycomb codes [15] but only graph-like edges for the CSS and X3Z3 Floquet codes [11]. As we explain below, these hyperedges degrade the MWPM decoder performance. As the noise bias increases, the thresholds for both honeycomb codes increase only by $\leq 6\times 10^{-4}$, where the thresholds at the infinite bias are only about $1.03-1.08$ of their thresholds at the depolarising noise.
@@
We find that there are differences, albeit modest ones, between the performance of the XYZ2 and P6 honeycomb codes under biased noise; these differences have not been pointed out before in the literature. While the XYZ2 honeycomb code has a better performance than the P6 honeycomb code for the code-capacity noise model, surprisingly we find that its performance is more inferior to that of the P6 Floquet code for the SDEM3 noise model, considering the similarity of the former to the bias-tailored XYZ2 static code [40]. However, owing to stabilisers being measured in two measurement subrounds, the XYZ2 Floquet code no longer possesses all the symmetries of the static version making the dynamical code not be able to inherit all the benefits of its static counterpart.
@@
The performance difference between the XYZ2 and P6 honeycomb codes can be understood from the distribution of hyperedge-like syndromes (namely, those with four triggered detectors) in the detector hypergraphs under the biased SDEM3 noise model. These hyperedges generally degrade the code’s performance when using a matching decoder since they must be decomposed (see Ref. [15] for a detailed description of the decoding in a hyperbolic version of the honeycomb code). In an infinitely biased SDEM3 noise model that includes two-qubit noise but no measurement errors, the P6 code has hyperedges resulting from only $Z$ errors occurring after the red check measurement subrounds, while the XYZ2 code has hyperedges resulting from certain $Z$ errors in all subrounds. There is no difference in the overall number of hyperedges in the two hypergraphs. However, one result of this different arrangement of the hyperedges is that two-qubit $ZZ$ errors can lead to hyperedge syndromes only in the XYZ2 Floquet code. This is explained in Appendix E. As a result of this, in an infintely biased regime, where Pauli errors after MPP gates are evenly distributed between $Z_{1}$, $Z_{2}$ and $Z_{1}Z_{2}$, there are more hyperedge syndromes triggered in the XYZ2 code than in the P6 code. To support this conclusion, we can see from Fig. 6(a) that the difference in thresholds between these codes is marginal under the code-capacity noise. Because honeycomb codes are prone to hyperedges error, better performance of the honeycomb codes might be expected using a correlated MWPM decoder such as the one used in Ref. [7].
@@
As a result of the shift of the thresholds with the noise bias, the subthreshold performance of the codes changes accordingly. As shown in Figs. 6(c) and 6(d), the subthreshold logical failure probability improves significantly for the X3Z3 code, while it becomes only slightly better for the honeycomb codes and deteriorates for the CSS code.
@@
Besides the threshold, another quantity that is of interest is the scaling of the code sub-threshold performance with the code distance. To show this, we plot the sub-threshold logical failure probability as a function of $d$ or $d_{\mathrm{eff}}$ for varying levels of noise bias in Fig. 7. The logical failure probabilities for all codes decrease exponentially with the code distance, i.e., $p_{L}\propto\mathrm{exp}(-\gamma d)$ or $p_{L}\propto\mathrm{exp}(-\gamma d_{\mathrm{eff}})$. Among all codes presented, the X3Z3 code has the largest logical error suppression rate $\gamma$. At higher noise bias, this error suppression rate becomes significantly larger for the X3Z3 code, moderately increases for the two honeycomb codes and decreases for the CSS Floquet code. The reason is that as the noise bias increases, a fixed subthreshold physical error rate moves relatively with respect to the shifting threshold, so that it becomes much further below the threshold for the X3Z3 Floquet code, moves moderately away from the threshold for the two honeycomb codes and becomes closer to the threshold for the CSS code [see Figs. 6(a) and 6(b)].
@@
The error suppression rate $\gamma$ is related to the error suppression factor $\Lambda$ proposed in Ref. [7] which is defined as the reduction factor in the logical failure probability as the code distance increases by 2. Mathematically, it is given by $\Lambda\equiv p_{L}(d)/p_{L}(d+2)=\mathrm{exp}(2\gamma)$ for the code-capacity noise and and similarly for the SDEM3 noise model but with $d$ replaced by $d_{\mathrm{eff}}$. In Table 1 of Appendix C.2 we list the values of $\Lambda$ corresponding to the $\gamma$ values shown in Fig. 7.
@@ VI No-go theorems for 1D decoding graphs of dynamical codes
Figure 8: An illustration of Theorems 1 (a) and 2 (b). We display examples of $Z$-detector graphs (edges correspond to $Z$ errors and vertices to detectors) that are forbidden by these theorems, for HDCs (a) and GDCs (b).
@@
Let $V$ be a set of detectors for a dynamical code. For every qubit and every time step (i.e., every fault location in the code) $(q,t)$, define a hyperedge $e_{(q,t)}=(v_{1},v_{2},\ldots)$, where the $v_{i}\in V$ are detectors that return $-1$ outcomes if a single $Z$ error at fault location $(q,t)$ occurs. The $\mathbf{Z}$-detector hypergraph (ZDH) is defined as $G=(V,E)$, where $E=\bigcup_{q,t}e_{(q,t)}$ is the set of all hyperedges. A $Z$-detector graph (ZDG) is a $Z$-detector hypergraph in which all hyperedges are edges (their size is 2).
@@
A honeycomb dynamical code (HDC), with “duration” $T$ and “detection delay” $k$, is a finite-depth measurement circuit acting on qubits of the honeycomb lattice (without boundary). The sets of measurements in the circuit, $\mathcal{M}_{t}$ ($t\in\{1,\ldots,T\}$), are composed of two-body Pauli measurements along coloured edges (either $r$, $g$, or $b$, for each time step) of the lattice, performed sequentially on a state stabilised by some group $\mathcal{S}$. The HDC obeys the following properties:
@@
    Detectors (for time steps $t>1$) are associated with plaquettes in the lattice (they have support only on qubits around their associated plaquette),
@@
    All single-qubit and non-trivial two-qubit errors occurring in time steps $1<t<T-k$ are detectable (a trivial error does not change the evolution of the system), and have syndromes of weight $>1$.
@@
Let us first comment on our definition of an HDC. The chosen properties are very natural. The non-commutativity of consecutive measurement subrounds ensures “local reversibility” of the code, so as to preserve the information  [55] (although note that we do not require the code to have logical qubits in our definition). The requirement that consecutive sets of measurements are supported on different edges is not very restrictive: if two (anti-commuting) measurements act on the same edge in consecutive time steps, we can replace the second with a Clifford gate and commute that to the end of the circuit. This merely changes the bases of subsequent measurements without changing their (anti-)commutation. As a result, we replace the HDC with $T$ time steps with one with $T^{\prime}<T$ time steps. We subtract a constant $k$ from the final time step $T$ in point (c) in order to avoid complications due to the final-time boundary, where errors close to this boundary can be detected by “reading out” logical qubits (in a memory experiment). That is, we consider errors occurring in “detection cells” [11] that are completed before time step $T$.
@@
A general dynamical code (GDC) is a finite-depth measurement circuit acting on qubits associated with vertices of a graph $G$. The sets of measurements $\mathcal{M}_{t}$ are two-body Pauli measurements associated with edges of $G$ (each qubit is in the support of one measurement in each time step), performed sequentially on a state stabilised by some group $\mathcal{S}$. A GDC obeys properties (a) and (c) of an HDC.
We will first show that an HDC cannot have a 1D-decodable ZDG and therefore its decoding cannot be equivalent to that of a collection of repetition codes. Note that a repetition code graph not only is 1D-decodable but also has a maximum degree of 2, which means there are no double edges between neighbouring vertices. Here, however, we allow for these double edges in our definition of 1D-decodability, and show that this more general property is also impossible. Such double edges naturally arise in Floquet codes, where they correspond to two-qubit undetectable errors (e.g., edge operators about to be measured).
Using the properties of the HDC as defined above, we now prove the first main theorem of our paper
@@
A honeycomb dynamical code with duration more than $43k+$ 4 (for detection delay $k$) cannot have a 1D-decodable $Z$-detector graph.
Proof.
Suppose there are $T$ time steps and $n$ qubits in the HDC. Note that we take an error occurring in time step $t$ to mean that it occurs after the measurements in $\mathcal{M}_{t}$. We will first restrict the ZDG: let it contain vertices corresponding to all detectors formed in time steps $t>1$, but only those edges corresponding to detectable errors, i.e., those occurring in time steps $1<t<T-k$. The number of vertices in this graph is at most $n(T-1)/6$, since in each time step no detectors can be formed on neighbouring plaquettes (see Lemma 1 in Appendix D), and we can select at most $n/6$ non-neighbouring plaquettes on the honeycomb lattice without boundary (corresponding to all plaquettes of a single colour). Meanwhile, $Z$ errors generate $n$ edges between each measurement subround, resulting in a total of $n(T-2-k)$ edges in the ZDG.
Consider the set $E_{t}$ of edges corresponding to single-qubit $Z$ errors from time step $t$. The maximum number of edges from $E_{t}$ between the same two vertices is 2. To see this, consider a scenario in which we have three or more edges between two vertices. Any pair of them corresponds to a weight-2 undetectable error, since it triggers no detectors. Consider three such undetectable errors: $Z_{i}Z_{j}$, $Z_{j}Z_{l}$ and $Z_{i}Z_{l}$. Since $i$, $j$ and $l$ are distinct qubits, one of these errors must be acting on non-adjacent qubits in the lattice, which must be detectable (see Lemma 2 in Appendix D).
Let us now create a bipartition of the qubits into sets $R$ and $S$ such that no qubits in $R$ are adjacent to one another in the honeycomb lattice, and similarly for qubits in $S$. This is possible since the honeycomb lattice is bipartite. Let us consider the subgraph of the ZDG that contains only edges corresponding to detectable $Z$ errors on qubits in $R$, ZDGR. ZDGR contains $n(T-2-k)/2$ edges. Suppose two vertices are connected in this graph by more than one edge. There is therefore at least one undetectable weight-2 error with support only on qubits in $R$. These two errors cannot occur in the same time step, owing to the bipartition of qubits (Lemma 2). Meanwhile, if the errors occur on different qubits, there is once again a non-trivial, undetectable error (they will anti-commute with some future measurements). Hence, they must occur on the same qubit, $i$, in different time steps. Since the pair of errors is undetectable, the measurements in between them must commute with $Z_{i}$. Hence, there can only be one such measurement, because consecutive measurements with support on $i$ anti-commute. Therefore, any two of these errors are separated by one time step and so there can only be two such errors. That is, vertices in ZDGR are only connected by at most two edges. Therefore, there is a spanning subgraph of ZDGR with at least $\lceil n(T-2-k)/4\rceil$ edges with no vertices in the subgraph connected by more than one edge.
Therefore, the whole ZDG has a subgraph with $\lceil n(T-2-k)/4\rceil$ edges and $\leq n(T-1)/6$ vertices such that no two vertices are connected by more than one edge. That is, there are many more edges than vertices in this subgraph (recall $T>3k+4$). Such a graph cannot have all vertices with neighbourhoods of size $\leq 2$. The size of the neighbourhood of some vertex in the full ZDG is therefore also greater than 2. ∎
This precludes the possibility of an HDC having a collection of repetition code graphs as its ZDG. Therefore, the route towards a large threshold under biased noise is impossible for HDCs. Indeed, for large $T$, there will be at least approximately $nT/(nT/6)=6$ times as many edges as vertices in the ZDG, and there is a subgraph with $3/2$ as many edges as vertices without any pair of vertices sharing more than one edge. This means the graph must have many cycles with length at least three edges.
It is clear that the codes examined in this paper obey the properties of an HDC. In particular, any Floquet codes related to the CSS Floquet code by Hadamard gates applied to a subset of the qubits (as is the case for the X3Z3 Floquet code) do not have 1D-decodable ZDGs. The reason we have focused on deformation by Hadamard gates is that $Z$ errors remain associated with graph-like edges in the deformed code’s ZDG, rather than hyperedges. While improved performance is possible for infinitely biased noise that produces hyperedges [32], we do not believe this would be the case for Clifford-deformed CSS Floquet codes owing to the high connectivity of the detector graphs of all Hadamard-deformed versions of the code.
Having defined GDCs, we now prove the second main theorem of our paper:
Theorem 2.
@@
A GDC with detection delay $k$ and duration greater than $22k+$ 2 (i.e., a constant) cannot have a $Z$-detector graph equivalent to a collection of disjoint repetition codes.
Proof.
There are $n/2$ measurements in a given time step. Regardless of how these measurements form detectors, there can be at most $n/2$ independent such detectors formed. However, there will in fact be far fewer, because each individual measurement anti-commutes with a measurement from the previous time step and so is random. Therefore, a detector must be formed from at least two measurements within a time step. We assume the GDC has a $Z$-detector graph. Each $Z$-error occurring in time steps $1<t<T-k$ is associated with an edge in the ZDG. Focusing on time steps $t>1$, there are therefore $<n(T-1)/2$ vertices in the ZDG and $n(T-2-k)$ edges. With a collection of repetition code graphs, $V\geq E$, for vertex (edge) number $V$ ($E$). Hence, $T>2k+2$ implies that the ZDG is not a collection of repetition code graphs. ∎

### v2

@@ V Results and Discussion
Figure 6: Thresholds ($p_{\mathrm{th}}$) of different Floquet codes as a function of noise bias $\eta$. Codes studied are X3Z3 (blue), CSS (green), P6 (red) and XYZ2 (orange). Results are computed for two different noise models: (a) code-capacity and (b) SDEM3. (a) Inset: Zoom-in threshold plots for the P6 (red) and XYZ2 (orange) codes. For better visualisation, we fit all curves with quadratic splines.
@@
Figure 7: Sub-threshold logical failure probability $p_{L}$ of different Floquet codes as a function of noise bias $\eta$. Codes studied are X3Z3 (blue), CSS (green), P6 (red) and XYZ2 (orange). Results are computed for two different noise models: (a) code-capacity and (b) SDEM3. Inset: Zoom-in subthreshold performance plots of the P6 (red) and XYZ2 (orange) codes. Results are calculated using (a) lattice size $20\times 30$ ($d=20$) for 30 QEC rounds with $p=0.72\%$ and (b) lattice size $24\times 36$ ($d_{\mathrm{eff}}=12$) for 36 QEC rounds with $p=0.5\%$. Each data point is averaged over $10^{6}-10^{8}$ Monte Carlo shots. For better visualisation, we fit all curves with quadratic splines.
@@
Figure 6 shows the thresholds of all codes for various levels of noise bias $\eta$ calculated for (a) code-capacity and (b) SDEM3 noise models. Each threshold is obtained from the intersection of the logical failure probabilities $p_{L}$ vs physical error rate $p$ curves of different code distances $d_{\mathrm{eff}}=d/2=6,8,10$ and 12. The data of $p_{L}$ as a function of $p$ for each code and various noise-bias strengths, used to obtain the thresholds, are presented in Figs. 12-15 of Appendix C.
@@
Since the code-capacity noise model, which considers only single-qubit noise, is a more benign model than the SDEM3 noise model, which also includes two-qubit and measurement errors, the code performance calculated under the code-capacity noise is better than that of the SDEM3 noise. For both noise models, as shown in Figs. 6(a) and 6(b), the performance of the X3Z3 code becomes increasingly better than those of all other tested codes as the noise bias increases. In particular, as the noise changes from fully depolarising to a pure dephasing type, the X3Z3 Floquet code’s threshold increases from $\approx 1.13\%$ to $\approx 3.09\%$ for the code-capacity noise and from $\approx 0.76\%$ to $\approx 1.08\%$ for the SDEM3 noise model. The threshold therefore increases by a factor of 2.7 and 1.4 for the code capacity and SDEM3 noise models, respectively.
It is interesting to note that for the fully depolarising code-capacity noise, all codes have the same threshold. This is because at every measurement subround under this noise model, all the above codes have two types of errors at each fault location (occurring with a total probability $32p/$ 3) that give rise to edges in the detector graph and one type (occurring with probability $p/3$) that produces hyperedges. For the CSS and X3Z3 codes, these hyperedges result from $Y$ Pauli errors. On the other hand, the hyperedge errors in the P6 and XYZ2 codes are those that anti-commute with the check measurements that occur just before and immediately after the errors [12, 22]. Therefore, the Pauli type of the hyperedge errors for the honeycomb codes varies between measurement subrounds. For example, $Z$ errors create hyperedges only when they occur between the $XX$ and $YY$ checks. In summary, all the above codes have the same performance for fully depolarising code-capacity noise because their (weighted) detector hypergraphs are all equivalent.
As shown in Figs. 6(a) and 6(b), while the CSS Floquet code has the same threshold as the X3Z3 code when the noise is fully depolarising, its threshold decreases with increasing noise bias, where its value is $0.752\%$ and $0.668\%$ at infinite bias for the code-capacity and SDEM3 noise, respectively. This decrease is due to the fact that the CSS Floquet code has pure $X$ and pure $Z$ detectors where in the presence of biased noise, half of the detectors, i.e., those with the same type of the dominant error, will become less useful in detecting the biased errors.
@@
Figures 6(a) and 6(b) show that the thresholds for the honeycomb codes have only minor improvements as the noise bias increases, where the thresholds for both honeycomb codes increase only by $\leq 6\times 10^{-4}$. Specifically, the thresholds of the honeycomb codes at $\eta=\infty$ are only about $1.03-1.08$ times larger than their thresholds at $\eta=0.5$ (where their fully-depolarising noise thresholds are $1.13\%$ and $0.585\%$ for the code-capacity and SDEM3 models, respectively). We note that for the SDEM3 noise, the thresholds of both honeycomb codes are lower than those of the X3Z3 and CSS Floquet codes at all noise biases. This is partly explained by noting that the SDEM3 noise model contains measurement errors which give rise to hyperedges in the decoding hypergraphs of the honeycomb codes [12] but only graph-like edges for the CSS and X3Z3 Floquet codes [13]. As we explain below, these hyperedges degrade the MWPM decoder performance.
@@
We find that there are differences, albeit modest ones, between the performance of the XYZ2 and P6 honeycomb codes under biased noise; these differences have not been pointed out before in the literature. While the XYZ2 honeycomb code has a better performance than the P6 honeycomb code for the code-capacity noise model, we find that its performance is inferior to that of the P6 Floquet code for the SDEM3 noise model. This is surprising, considering the similarity of the former to the bias-tailored XYZ2 static code [46]. However, owing to stabilisers being measured in two measurement subrounds, the XYZ2 Floquet code no longer possesses all the symmetries of the static version making the dynamical code not able to inherit all the benefits of its static counterpart.
@@
The performance difference between the XYZ2 and P6 honeycomb codes can be understood from the distribution of hyperedge-like syndromes (namely, those with four triggered detectors) in the detector hypergraphs under the biased SDEM3 noise model. These hyperedges generally degrade the code’s performance when using a matching decoder since they must be decomposed into edges (see Ref. [12] for a detailed description of the decoding in a hyperbolic version of the honeycomb code). An infinitely biased SDEM3 noise model that includes Pauli noise but no measurement errors can already form hyperedges in the honeycomb codes’ detector hypergraphs. In particular, the P6 code can have hyperedges resulting from $Z$ errors occurring only after the red check measurement subrounds, while the hyperedges in XYZ2 code can result from certain $Z$ errors happening in any subround. While there is no difference in the overall number of hyperedges resulting from single-qubit $Z$ errors in the two hypergraphs, the different arrangement of the hyperedges due to the single-qubit errors in these two codes gives rise to different kinds of syndromes for the two-qubit $ZZ$ errors. Specifically, $ZZ$ errors can lead to hyperedge syndromes in the XYZ2 Floquet code but only edge-like syndromes for the P6 code (see Appendix D for details). As a result, in an infinitely biased regime, where Pauli errors after MPP gates are evenly distributed between $Z_{1}$, $Z_{2}$ and $Z_{1}Z_{2}$, there are more hyperedge syndromes triggered in the XYZ2 code than in the P6 code. This is the reason underlying the difference in the thresholds, as shown in Fig. 6(b), of the two honeycomb codes under the SDEM3 noise. Since honeycomb codes are prone to hyperedge errors, better performance of the honeycomb codes might be expected using a correlated MWPM decoder such as the one used in Ref. [8].
@@
As a result of the shift of the thresholds with the noise bias, the subthreshold performance of the codes changes accordingly. As shown in Figs. 7(a) and 7(b), the subthreshold logical failure probability improves significantly for the X3Z3 code, while it becomes only slightly better for the honeycomb codes and deteriorates for the CSS code.
Figure 8: Exponential suppression of the sub-threshold logical failure probabilities $p_{L}$ with respect to $d$ or $d_{\mathrm{eff}}$ for different Floquet codes: CSS (green), XYZ2 (orange), P6 (red), and X3Z3 (blue). Results are calculated for different noise models: (a,b) code-capacity and (c,d) SDEM3 noise models, with a physical error rate $p=0.55\%$. They are also computed using different bias strengths: (a,c) $\eta=1$ and (b,d) $\eta=99$. All curves can be fitted to an exponential decay function $f\propto\mathrm{exp}(-\gamma d)$ or $f\propto\mathrm{exp}(-\gamma d_{\mathrm{eff}})$ where $\gamma$ depends on the bias strength $\eta$ and is an increasing function of $(p_{\mathrm{th}}-p)$. Each data point is averaged over $10^{5}-10^{9}$ shots.
@@
Besides thresholds, another quantity that is of important is the scaling of the code sub-threshold performance with the code distance. To show this, we plot the sub-threshold logical failure probability as a function of $d$ or $d_{\mathrm{eff}}$ for varying levels of noise bias in Fig. 8. As shown, the logical failure probabilities for all codes decrease exponentially with the code distance, i.e., $p_{L}\propto\mathrm{exp}(-\gamma d)$ or $p_{L}\propto\mathrm{exp}(-\gamma d_{\mathrm{eff}})$. Among all codes presented, the X3Z3 code has the largest logical error suppression rate $\gamma$. At higher noise bias, this error suppression rate becomes significantly larger for the X3Z3 code, moderately increases for the two honeycomb codes and decreases for the CSS Floquet code. The reason is that as the noise bias increases, a fixed subthreshold physical error rate moves relatively with respect to the shifting threshold, so that it becomes much further below the threshold for the X3Z3 Floquet code, moves moderately away from the threshold for the two honeycomb codes and becomes closer to the threshold of the CSS code [see Figs. 6(a) and 6(b)].
We also numerically simulate the performance of the X3Z3 Floquet code with a twisted periodic boundary condition. This twisted code has a pure $Z$-type logical operator with a length that scales quadratically with the code distance of the untwisted code. As a result, we find better thresholds and more favorable subthreshold performance for the twisted code over the untwisted one. While these performance benefits hold for the infinite-bias code-capacity noise, they do not apply to other noise biases and models tested (See Appendix E for details).
@@
The error suppression rate $\gamma$ is related to the error suppression factor $\Lambda$ introduced in Ref. [66] which is defined as the reduction factor in the logical failure probability as the code distance increases by 2. Mathematically, it is given by $\Lambda\equiv p_{L}(d)/p_{L}(d+2)=\mathrm{exp}(2\gamma)$ for the code-capacity noise and similarly for the SDEM3 noise model but with $d$ replaced by $d_{\mathrm{eff}}$. In Table 1 of Appendix C.2 we list the values of $\Lambda$ corresponding to the $\gamma$ values shown in Fig. 8.
@@ VI No-go theorems for 1D decoding graphs of dynamical codes
Figure 9: An illustration of (a) Theorem 1 and (b) Theorem 2. We display examples of the $Z$-detector graphs (edges correspond to $Z$ errors and vertices to detectors) that are forbidden by these theorems, for (a) HDCs and (b) GDCs.
@@
Let $V$ be a set of detectors for a dynamical code. For every qubit and every time step $(q,t)$, i.e., every fault location in the code, define a hyperedge $e_{(q,t)}=(v_{1},v_{2},\ldots)$, where the $v_{i}\in V$ are detectors that return $-1$ outcomes if a single $Z$ error occurs at the fault location $(q,t)$. The $\mathbf{Z}$-detector hypergraph (ZDH) is defined as $G=(V,E)$, where $E=\bigcup_{q,t}e_{(q,t)}$ is the set of all hyperedges. A $Z$-detector graph (ZDG) is a $Z$-detector hypergraph in which all hyperedges are edges (their size is 2).
@@
A honeycomb dynamical code (HDC), with “duration” $T$ and constant “initial(final)-time boundary offsets” $k_{i}$ ($k_{f}$), is a finite-depth measurement circuit acting on qubits of the honeycomb lattice (without boundary). The sets of measurements in the circuit, $\mathcal{M}_{t}$ ($t\in\{1,\ldots,T\}$), are composed of two-body Pauli measurements along coloured edges (either $r$, $g$, or $b$, for each time step) of the lattice, performed sequentially on a state stabilised by some group $\mathcal{S}$. The HDC obeys the following properties:
@@
    Detectors (for time steps $t>k_{i}$) are associated with plaquettes in the lattice (they have support only on qubits around their associated plaquette),
@@
    If a single-qubit or two-qubit error occurring in time steps $k_{i}<t<T-k_{f}$ anti-commutes with future measurements, it is detectable, unless the two-qubit error is the same as an edge operator just measured or to be measured in the next time step.
  4. (d)
@@
    All single-qubit errors in time steps $k_{i}<t<T-k_{f}$ have syndromes of weight $>1$.
@@
Let us first comment on our definition of an HDC. The chosen properties are very natural. The anti-commutation of consecutive measurement subrounds ensures “local reversibility” of the code, so as to preserve the quantum information and its locality  [67] (although note that we do not require the code to have logical qubits in our definition). The requirement that consecutive sets of measurements are supported on different edges is not very restrictive: if two (anti-commuting) measurements act on the same edge in consecutive time steps, we can replace the second with a Clifford gate and commute that to the end of the circuit. This merely changes the bases of subsequent measurements without changing their (anti-)commutation. Since the Clifford gates are unimportant, we can ignore them and hence we end up with a shorter duration $T$. We restrict our attention to errors occurring in the temporal “bulk” of the code, $k_{i}<t<T-k_{f}$. This avoids complications due to errors occurring close to the initial and final-time boundaries where detectors here may be formed in different ways than the bulk detectors. Specifically, the detectors associated with errors at the initial and final time boundaries are obtained by comparing the edge or plaquette measurement values with the initial state and final read-out of the physical qubits, respectively. The exclusion of these final time boundary detectors is also required here since we do not include final read-out measurements in our definition of an HDC. In summary, we consider errors occurring in all “detection cells” [13] that are completed before time step $T$ and begin at time step $t\geq 1$, where a detection cell consists of all the spacetime points at which a detector can identify errors.
We will show that an HDC cannot even have a 1D-decodable ZDG, which implies that its decoding graph cannot be equivalent to a set of the much simpler repetition code’s graph. Note that a repetition code graph is not only 1D-decodable but also has a maximum degree of 2, which means there are no double edges between neighbouring vertices. Here, however, we allow for these double edges in our definition of 1D-decodability, and show that this more general property is also impossible. Such double edges naturally arise in Floquet codes, where they correspond to two-qubit undetectable errors (i.e., edge operators just measured or about to be measured).
Using the properties of the HDC as defined above, we now present the first main theorem of our paper (see Appendix F for a proof):
Theorem 1.
A honeycomb dynamical code with duration more than $33k_{f}+k_{i}+$ 3 cannot have a 1D-decodable $Z$-detector graph, where $k_{i}$ and $k_{f}$ are the initial and final time-boundary offsets.
The above theorem implies that no matter how an HDC is bias-tailored, it likely can never achieve as high a threshold as the repetition code or the bias-tailored stabiliser codes. Indeed, for large $T$ ($T\gg k_{i},k_{f}$), there will be at least approximately $6$ times as many edges as vertices in the ZDG, and there is a subgraph with $3/2$ as many edges as vertices without any pair of vertices sharing more than one edge (see Appendix F). This means the graph must have many cycles with length at least three edges.
It is clear that the codes examined in this paper obey the properties of an HDC. In particular, any Floquet codes related to the CSS Floquet code by Hadamard gates applied to a subset of the qubits (as is the case for the X3Z3 Floquet code) do not have 1D-decodable ZDGs. The reason we have focused on dynamical codes that have ZDGs, e.g., Hadamard-deformed CSS codes, instead of codes with hyperedges in their $Z$-detector hypergraphs, is because their decoding graphs are immediate generalisations of that of the CSS code. The fact that 1D-decodable ZDGs are not possible for an HDC suggests that there are not enough independent symmetries to reduce its decoding graph to the extent possible for stabiliser codes. It is interesting to note that even with hyperedges in the $Z$-detector hypergraph, a stabiliser code can still be bias-tailored [38] such that it possesses large enough symmetries to render its thresholds close to the hashing bound [68, 69]. While such high-thresholds are possible for static codes, we do not believe this would be the case for HDCs.
@@
A general dynamical code (GDC) is a finite-depth measurement circuit acting on qubits associated with vertices of a graph $G$. The sets of measurements $\mathcal{M}_{t}$ are two-body Pauli measurements associated with edges of $G$ (each qubit is in the support of one measurement in each time step), performed sequentially on a state stabilised by some group $\mathcal{S}$. A GDC obeys properties (a) and (d) of an HDC.
Having defined GDCs, we now provide the second main theorem of our paper:
@@
A GDC with a duration at least $22k_{f}+k_{i}+$ 2 (i.e., a constant), where $k_{i}$ and $k_{f}$ are the initial and final time-boundary offsets, respectively, cannot have a $Z$-detector graph equivalent to a collection of disjoint repetition codes.